# Solving the Cold Start Problem on One’s Own as an End User via Preference Transfer

### v1

@@ 3 Pretender
We propose Pretender (PREference Transfer by END usERs) to solve the preference transfer problem. We first describe the gerneral framework and then provide variants for specific settings.
@@ 3.1 Formulating the Problem as Distance Minimization
[…]
@@
Although we have formulated the problem as distance minimization between distributions, this problem remains challenging because selection of items is combinatorial. This is in contrast to other minimization problems, where optimization is performed over weights and/or coordinates of data points, making the problem continuous and often convex. However, end users cannot “thumbs up 0.2 points” or alter the features of the items in the service, meaning that we cannot sidesitep the combinatorial nature of the problem. We address this challenge in the following.
@@ 3.2 Optimization
    (Continuous Optimization) Optimize the weights ${\bm{w}}\in\left[0,\frac{1}{K}\right]^{2m}\cap\Delta_{2m}$ of the items in the target service such that the weighted empirical distribution $\mu_{T}^{\bm{w}}=\sum_{j=1}^{2m}{\bm{w}}_{i}\delta_{({\bm{x}}_{i_{j}},y_{j})}$ is close to the source distribution $\mu_{S}$, where $\Delta_{d}$ denotes the $(d-1)$-dimensional probability simplex. This is achieved by solving the following optimization problem:
@@
where (a) follows because the weight $(1/K,\ldots,1/K)$ of the empirical meassure is in the feasible set. Therefore, the continuous solution is at least as good as the combinatorial solution. However, we need to carefully round the solution to obtain the final output to ensure that the rounding process does not degrade the quality of the solution much. This is the main challenge from a theoretical perspective.
Random Selection. We employ a randomized approach. We first point out that the distribution Bernoulli$(K{\bm{w}}_{j})$ is well-defined as we set the optimization domain to ${\bm{w}}\in\left[0,\frac{1}{K}\right]^{2m}\cap\Delta_{2m}$ and $0\leq K{\bm{w}}_{j}\leq 1$ holds. Let $\tilde{{\bm{w}}}\in\{0,1/K\}^{2m}$ be the sample weights after the random selection, i.e., $\tilde{{\bm{w}}}_{j}=\frac{1}{K}I_{j}$. Then,
@@
$\text{Var}\left[\sum_{j=1}^{2m}I_{j}\right]$ =$\stackrel{{\scriptstyle\text{(a)}}}{{=}}\sum_{j=1}^{2m}\text{Var}[I_{j}]\stackrel{{\scriptstyle\text{(b)}}}{{=}}\sum_{j=1}^{2m}K{\bm{w}}_{j}(1-K{\bm{w}}_{j})\stackrel{{\scriptstyle\text{(c)}}}{{\leq}}\sum_{j=1}^{2m}K{\bm{w}}_{j}\stackrel{{\scriptstyle\text{(d)}}}{{=}}K,$ (14)
where (a) follows from the independence of the Bernoulli random variables, (b) follows from the variance of the Bernoulli random variables, (c) follows from ${\bm{w}}_{j}\geq 0$, and (d) follows from $\sum_{j=1}^{2m}{\bm{w}}_{j}=1$. Therefore, the standard deviation is the order of $\sqrt{K}$, and the number of selected items is concentrated around $[K-O(\sqrt{K}),K+O(\sqrt{K})]$. This implies that the selection process does not blow up the error too much. We will elaborate this discussion in the following sections.
@@ 3.3 Pretender for MMD
$={\bm{w}}^{\top}{\bm{K}}^{TT}{\bm{w}}-\frac{2}{n}\mathbb{1}^{\top}{\bm{K}}^{ST}{\bm{w}}+\text{const.},$. , (21)
@@
Let $\tilde{{\bm{w}}}^{\prime}_{j}=\frac{1}{K}$ if the $j$-th item is in the final output $\mathcal{D}_{T}$ and 00 otherwise. We have
@@
Under assumption 3.2, for any $\epsilon>0$ and $K\in\mathbb{Z}_{D}$, there exists $L$ and $R$ such that by running the Frank-Wolfe algorithm with $L$ iterations and repeating the rounding process $R$ times and choose the best $\mathcal{D}^{*}_{T}$ with smallest $\text{MMD}(\mu_{T}^{\mathcal{D}^{*}_{T}},\mu_{S})$, we have
@@
with probability at least $01-2\delta>$ 0 because $\delta<\frac{1}{2}$. Repeating the rounding process $R=\Theta((1-2\delta)^{-1})$ times, we can obtain the final output with the desired guarantee. ∎
@@
This corollary shows that if the target service trains the model sufficiently well so that the training error $\frac{1}{|\mathcal{D}_{T}|}\sum_{({\bm{x}},y)\in\mathcal{D}_{T}}\ell_{T}(f_{T}({\bm{x}};\theta),y)$ is small and if we prepare a sufficient number of source items and click sufficiently many items following Predenter, we can make the right hand side of Eq. 33 arbitrarily small and we can ensure that the trained model incurs a small error on the source items, meaning that it effectively captures the user’s preferences from the source service.
@@ 3.4 Pretender for the Wasserstein Distance
Each element $x_{j}$ lies is the unit cube $[0,1]^{d}$ and the cost function $\|x-x^{\prime}\|$ is the Euclidean distance.
@@
Let $\mathcal{F}_{\text{Lip}}$ be the set of $1$-Lipschitz functions with $f(\frac{\mathbb{1}}{2})=0$. $\mathcal{F}_{\text{Lip}}$ can be $L_{\infty}$-covered by ~$\tilde{\mathcal{F}}_{\varepsilon}$, i.e., for any $f\in\mathcal{F}_{\text{Lip}}$, there exists $\tilde{f}\in\tilde{\mathcal{F}}_{\varepsilon}$ such that $\|f-\tilde{f}\|_{\infty}\leq\varepsilon$, with $|f(x)|\leq\frac{\sqrt{d}}{2},\forall f\in\tilde{\mathcal{F}}_{\varepsilon}$ and |~$|\tilde{\mathcal{F}}_{\varepsilon}|\leq\exp\left(\log(3)\left(\frac{2\sqrt{d}}{\varepsilon}\right)^{d}\right)$.
@@
Let $\tilde{{\bm{w}}}^{\prime}_{j}=\frac{1}{K}$ if the $j$-th item is in the final output $\mathcal{D}_{T}$ and 00 otherwise. For any $\varepsilon>0$,
@@
We set $\varepsilon=\sqrt{d}\left(2^{\frac{d-2}{d+2}}\log^{\frac{1}{d+2}}(3)K^{-\frac{1}{d+2}}+\frac{1}{2}\sqrt{\frac{1}{K}\log\frac{1}{\delta}}\right)$ in Lemma 3.12. We bound |~$|\tilde{\mathcal{F}}_{\varepsilon}|$, set $\delta\leftarrow\delta/|\tilde{\mathcal{F}}_{\varepsilon}|$ in Proposition 3.14, and apply the union bound. We then apply Proposition 3.13, Lemma 3.14, Lemma 3.12, and the Kantrovich-Rubinstein duality to obtain the desired result. The full proof is in Appendix H. ∎
@@ 4.1 Discussion on Variants
We also note that our analysis can be extended to other metrics as longs as the continuous optimization problem (Eq. 7) is tractable and the convering number of the test function class $\mathcal{F}$ is bounded by the same argument of Section 3.4. For example, our analysis can be applied to the discrepancy distance [30], yielding a sharp generalization bound when the model is linear and the loss is the squared loss. Practitioners can choose metrics that best suits their specific application.
@@ 4.2 Optimum Value is Not Monotone
We prove the proposition by a counterexample. Let $I_{S}=\{1\},I_{T}=\{1,2\},{\bm{x}}_{1}=0,{\bm{x}}_{2}=1$, and $\mathcal{D}_{S}=\{(1,1)\}$. When $K=1$, the optimum value is 00 by selecting $\mathcal{D}_{T}=\{(1,1)\}$. When $K=2$, the optimum value is non-zero because we need to select items other than $(1,1)$, and $\mathcal{D}_{S}\neq\mathcal{D}_{T}$. ∎
@@ 5.2 Quadrature
[…]
@@ 5.3 User-side Realization
User-side realization refers to the concept in which end users implement desired functionalities on their own without requiring modifications to the service itself. Many users experience dissatisfaction with services. Even if they want some functinoalities and request them to the service provider, the provider may not implement them due to various reasons such as cost, complexity, and simple negligence. After all, service providers are not volunteers but businesses. In such cases, the only options users have are not satisfactory, keep using the service despite their dissatisfaction or leave the service. User-side realization provides a proactive alternative to this dilemma. This concept has been explored in various fields such as recommender systems [43, 45, 47], search engines [44, 42, 11, 40, 36], and privacy [46]. The main advantage of the user-side realization is that it can be used in services that do not have special functionalities to address the problem, and it broadens the scope of the applicability of the solution. For a more comprehensive discussion on user-side realization, we refer the reader to the Ph.D. thesis by [48]. Our approach constitutes a novel application of user-side realization to the cold start problem in recommender systems.
@@ 6.1 Convergence Analysis
Experimental Setting. We use the MovieLens 100K dataset [17]. We consider rating $\geq 4$ as positive feedback (i.e., thumbs up) and rating $<4$ as negative feedback (i.e., thumbs down). We create two virtual services $S$ and $T$. We include each movie in $S$ with probability $0.5$. We carry out the same process independently to create $T$. As a result, these services share about half of the movies. We use genres (e.g., Action, Adventure, Animation) and the release year as the features of the movies. They are encoded as a $90$-dimensional multi-hot vector. We concatenate $Cy\in\{0,C\}$ to the feature vector to define the discrete distribution, where $y\in\{0,1\}$ indicates thumbs up or thumbs down, and $C=10$ is the hyperparameter that controls the emphasis on the label when we measure the distance between two data points. We focus on user 308, who is the first user who has $10$ thumbs up and $10$ thumbs down, and user 21, who is the first user who has more than $100$ ratings. For each user $u$, we aim to transfer the preference $\mathcal{D}_{u}\cap\mathcal{D}_{S}$ by choosing $K$ items from $\mathcal{D}_{T}$. We use MMD with the Gaussian kernel and bandwidth $\sigma=1$ as the distance measure. We use $L=1000$ iterations of the Frank-Wolfe algorithm and $R=100$ trials of the rounding process. We set $K=1,2,4,8,\ldots,128$ to see the convergence behavior.
@@ 6.2 Quantitative Evaluation
Datasets. We evaluate our method using datasets from three domains: MovieLens 100k (movies), Last.fm (music) [7], and Amazon-Home-and-Kitchen (e-commerce) [20, 32]. We use the same setting as the convergence analysis for the MovieLens 100K dataset except that we use all users. For the Last.fm dataset, we consider musicians with positive weights as thumbs up and we sample the same number of random musicians for thumbs down. We use the tags (e.g., metal, 80’s, chillout) as the features of the artists. As the number of tags is large, we apply the principal component analysis (PCA) to reduce the dimension to $50$ and standardize the variance. For the Amazon-Home-and-Kitchen dataset, we consider ratings $\geq 4$ as thumbs up and ratings $<4$ as thumbs down. We use the bag-of-words representation of the product reviews. We apply PCA to reduce the dimension to $50$ and standardize the variance. We consider two different settings for defining source and target services. The first setting, with intersection, is identical to the one used in the convergence analysis. The second setting, no intersection, is that we split the items into source and target services without any intersection. This setting is more challenging because the methods cannot copy any source items to the target service. For Last.fm and Amazon-Home-and-Kitchen datasets, we use the Gaussian kernel with the bandwidth $\sigma=10$ to reflect the scale of the features. We set $K=100$ for all datasets.
Baselines. Since there are no existing methods that solve the cold start problem by an end user, we use two simple baselines to validate that the proposed method is reasonablly effective. The first baseline is random selection, which randomly selects $K$ items from $\mathcal{D}_{T}$. The second baseline is the greedy selection, which selects $K$ items in $J_{T}$ that have the smallest distance $d(x,\mathcal{D}_{S})=\min_{x^{\prime}\in\mathcal{D}_{S}}\|x-x^{\prime}\|_{2}$ to the source items $\mathcal{D}_{S}$. Additionally, we report the performance of the optimal solution of the continuous optimization problem.
Results. Table 2 shows MMD values between the source preference and the output of the methods. Lower is better. Pretender is much better than the baseline methods and is close to the optimal continuous solution. Note that checking the closeness is useful in practice as well. The continuous solution is obtained as a byproduct of Pretender. After we obtain the final output, we can verify that the final solution is indeed a reasonable one just by comparing the two values. We observe that the greedy method is worse than the random method. We hypothesis that it is because the greedy method chooses items that are too similar to some source items and the diversity is lost. We also observe that Pretender is robust to the setting with no intersection. The performance is slightly worse than in the setting with intersection but is still reasonablly well.

### v2

@@ 1 Introduction
Reproducibility: Our code is available at https://github.com/joisino/pretender.
@@ 3 Pretender
We propose Pretender (PREference Transfer by END usERs) to solve the preference transfer problem. We first describe the general framework and then provide variants for specific settings.
@@ 3.1 Formulating the Problem as Distance Minimization
$(\text{loss on the source data})$ $=\frac{1}{|\mathcal{D}_{S}|}\sum_{(i,y_{i})\in\mathcal{D}_{S}}\ell_{T}(f_{T}({\bm{x}}_{i};\theta),y_{i})$ (3)
$=\int\ell_{T}(f_{T}({\bm{x}};\theta),y)\,d\mu_{S}({\bm{x}},y)$ (4)
$\leq\int\ell_{T}(f_{T}({\bm{x}};\theta),y)\,d\mu_{T}^{\mathcal{D}_{T}}({\bm{x}},y)+L\cdot W_{1}(\mu_{T}^{\mathcal{D}_{T}},\mu_{S})$ (5)
@@
Although we have formulated the problem as distance minimization between distributions, this problem remains challenging because selection of items is combinatorial. This is in contrast to other minimization problems, where optimization is performed over weights and/or coordinates of data points, making the problem continuous and often convex. However, end users cannot “thumbs up 0.2 points” or alter the features of the items in the service, meaning that we cannot sidestep the combinatorial nature of the problem. We address this challenge in the following.
@@ 3.2 Optimization
    (Continuous Optimization) Optimize the weights ${\bm{w}}\in\left[0,\frac{1}{K}\right]^{2m}\cap\Delta_{2m}$ of the items in the target service such that the weighted empirical distribution $\mu_{T}^{\bm{w}}=\sum_{j=1}^{2m}{\bm{w}}_{i}\delta_{({\bm{x}}_{i_{j}},y_{j})}$ is close to the source distribution $\mu_{S}$, where $\Delta_{d}$ denotes the (d1)(d-1)-dimensional probability simplex. This is achieved by solving the following optimization problem:
@@
where (a) follows because the weight $(1/K,\ldots,1/K)$ of the empirical measure is in the feasible set. Therefore, the continuous solution is at least as good as the combinatorial solution. However, we need to carefully round the solution to obtain the final output to ensure that the rounding process does not degrade the quality of the solution much. This is the main challenge from a theoretical perspective.
[…]
@@
$\text{Var}\left[\sum_{j=1}^{2m}I_{j}\right]$ $\stackrel{{\scriptstyle\text{(a)}}}{{=}}\sum_{j=1}^{2m}\text{Var}[I_{j}]\stackrel{{\scriptstyle\text{(b)}}}{{=}}\sum_{j=1}^{2m}K{\bm{w}}_{j}(1-K{\bm{w}}_{j})\stackrel{{\scriptstyle\text{(c)}}}{{\leq}}\sum_{j=1}^{2m}K{\bm{w}}_{j}\stackrel{{\scriptstyle\text{(d)}}}{{=}}K,$ (14)
where (a) follows from the independence of the Bernoulli random variables, (b) follows from the variance of the Bernoulli random variables, (c) follows from ${\bm{w}}_{j}\geq 0$, and (d) follows from $\sum_{j=1}^{2m}{\bm{w}}_{j}=1$. Therefore, the standard deviation is the order of $\sqrt{K}$, and the number of selected items is concentrated around $K-O(\sqrt{K}),K+O(\sqrt{K})]$. This implies that the selection process does not blow up the error too much. We will elaborate this discussion in the following sections.
@@ 3.3 Pretender for MMD
$={\bm{w}}^{\top}{\bm{K}}^{TT}{\bm{w}}-\frac{2}{n}\mathbb{1}^{\top}{\bm{K}}^{ST}{\bm{w}}+\text{const.},$ (21)
@@
Let $\tilde{{\bm{w}}}^{\prime}_{j}=\frac{1}{K}$ if the $j$-th item is in the final output $\mathcal{D}_{T}$ and 0 otherwise. We have
@@
Under assumption 3.2, for any $\epsilon>0$ and $K\in\mathbb{Z}_{+}$, there exists $L$ and $R$ such that by running the Frank-Wolfe algorithm with $L$ iterations and repeating the rounding process $R$ times and choose the best $\mathcal{D}^{*}_{T}$ with smallest $\text{MMD}(\mu_{T}^{\mathcal{D}^{*}_{T}},\mu_{S})$, we have
@@
with probability at least $1-2\delta>0$ because $\delta<\frac{1}{2}$. Repeating the rounding process $R=\Theta((1-2\delta)^{-1})$ times, we can obtain the final output with the desired guarantee. ∎
@@
This corollary shows that if the target service trains the model sufficiently well so that the training error $\frac{1}{|\mathcal{D}_{T}|}\sum_{({\bm{x}},y)\in\mathcal{D}_{T}}\ell_{T}(f_{T}({\bm{x}};\theta),y)$ is small and if we prepare a sufficient number of source items and click sufficiently many items following Pretender, we can make the right hand side of Eq. 33 arbitrarily small and we can ensure that the trained model incurs a small error on the source items, meaning that it effectively captures the user’s preferences from the source service.
@@ 3.4 Pretender for the Wasserstein Distance
Each element $x_{j}$ lies is the unit cube $0,1]^{d}$ and the cost function $\|x-x^{\prime}\|$ is the Euclidean distance.
@@
Let $\mathcal{F}_{\text{Lip}}$ be the set of $1$-Lipschitz functions with $f(\frac{\mathbb{1}}{2})=0$. $\mathcal{F}_{\text{Lip}}$ can be $L_{\infty}$-covered by $\tilde{\mathcal{F}}_{\varepsilon}$, i.e., for any $f\in\mathcal{F}_{\text{Lip}}$, there exists $\tilde{f}\in\tilde{\mathcal{F}}_{\varepsilon}$ such that $\|f-\tilde{f}\|_{\infty}\leq\varepsilon$, with $|f(x)|\leq\frac{\sqrt{d}}{2},\forall f\in\tilde{\mathcal{F}}_{\varepsilon}$ and $|\tilde{\mathcal{F}}_{\varepsilon}|\leq\exp\left(\log(3)\left(\frac{2\sqrt{d}}{\varepsilon}\right)^{d}\right)$.
@@
Let $\tilde{{\bm{w}}}^{\prime}_{j}=\frac{1}{K}$ if the $j$-th item is in the final output $\mathcal{D}_{T}$ and 0 otherwise. For any $\varepsilon>0$,
@@
We set $\varepsilon=\sqrt{d}\left(2^{\frac{d-2}{d+2}}\log^{\frac{1}{d+2}}(3)K^{-\frac{1}{d+2}}+\frac{1}{2}\sqrt{\frac{1}{K}\log\frac{1}{\delta}}\right)$ in Lemma 3.12. We bound $|\tilde{\mathcal{F}}_{\varepsilon}|$, set $\delta\leftarrow\delta/|\tilde{\mathcal{F}}_{\varepsilon}|$ in Proposition 3.14, and apply the union bound. We then apply Proposition 3.13, Lemma 3.14, Lemma 3.12, and the Kantrovich-Rubinstein duality to obtain the desired result. The full proof is in Appendix H. ∎
@@ 4.1 Discussion on Variants
We also note that our analysis can be extended to other metrics as long as the continuous optimization problem (Eq. 7) is tractable and the covering number of the test function class $\mathcal{F}$ is bounded by the same argument of Section 3.4. For example, our analysis can be applied to the discrepancy distance [30], yielding a sharp generalization bound when the model is linear and the loss is the squared loss. Practitioners can choose metrics that best suits their specific application.
@@ 4.2 Optimum Value is Not Monotone
We prove the proposition by a counterexample. Let $I_{S}=\{1\},I_{T}=\{1,2\},{\bm{x}}_{1}=0,{\bm{x}}_{2}=1$, and $\mathcal{D}_{S}=\{(1,1)\}$. When $K=1$, the optimum value is 0 by selecting $\mathcal{D}_{T}=\{(1,1)\}$. When $K=2$, the optimum value is non-zero because we need to select items other than $(1,1)$, and $\mathcal{D}_{S}\neq\mathcal{D}_{T}$. ∎
@@ 5.2 Quadrature
Quadrature is a technique to approximate the integral of a function by summing the function values at a finite number of points [41, 21, 18, 1, 19]. This is essentially equivalent to finding a discrete measure that approximates the given measure. Quadrature is widely used in numerical analysis and machine learning. One of the common applications is coresets [23, 50, 6, 33], which is a small set of training points that approximates the loss of the model on the entire training set. Our approach can be seen as a quadrature of $\int\ell_{T}\,d\mu_{S}$ with points $\mathcal{D}_{T}$. The main difference is that standard quadrature methods use arbitrary points and/or weights and sidestep combinatorial problems [2, 5], while our approach uses only the items in the target set, which naturally lead to the combinatorial optimization problem. For example, quadrature methods based on the Frank-Wolfe algorithm [58, 8, 2] output sparse weights, but they may choose the same items repeatedly in general, and the resulting weights are not uniform. Such output cannot be applied to our setting because users cannot thumbs up the same item multiple times in most services. Some approaches [23, 13] such as Kernel thinning [13] output a subset of the input points with uniform weights, which is similar to our approach. However, these methods assume that the candidate points are sampled according to the distribution being approximated. If this assumption does not hold, as in our case, these methods cannot be directly applied. Our proposed method can be used even when the input points are arbitrary. Other methods [56, 57] employ submodular optimization and greedy algorithms, achieving a $(1-\frac{1}{e})$-approximation ratio. However, the gap of $(1-\frac{1}{e})$ does not vanish as the number of items increases. By contrast, our approach can achieve the vanishing error as Corollary 3.10 shows thanks to the continuous optimization approach and the careful rounding process. To the best of our knowledge, our work is the first to provide a theoretical guarantee for such a general and combinatorial setting. We believe that this result is of independent interest.
@@ 5.3 User-side Realization
User-side realization refers to the concept in which end users implement desired functionalities on their own without requiring modifications to the service itself. Many users experience dissatisfaction with services. Even if they want some functionalities and request them to the service provider, the provider may not implement them due to various reasons such as cost, complexity, and simple negligence. After all, service providers are not volunteers but businesses. In such cases, the only options users have are not satisfactory, keep using the service despite their dissatisfaction or leave the service. User-side realization provides a proactive alternative to this dilemma. This concept has been explored in various fields such as recommender systems [43, 45, 47], search engines [44, 42, 11, 40, 36], and privacy [46]. The main advantage of the user-side realization is that it can be used in services that do not have special functionalities to address the problem, and it broadens the scope of the applicability of the solution. For a more comprehensive discussion on user-side realization, we refer the reader to the Ph.D. thesis by [48]. Our approach constitutes a novel application of user-side realization to the cold start problem in recommender systems.
@@ 6.1 Convergence Analysis
Experimental Setting. We use the MovieLens 100K dataset [17]. We consider rating $\geq 4$ as positive feedback (i.e., thumbs up) and rating <4<4 as negative feedback (i.e., thumbs down). We create two virtual services $S$ and $T$. We include each movie in $S$ with probability $0.5$. We carry out the same process independently to create $T$. As a result, these services share about half of the movies. We use genres (e.g., Action, Adventure, Animation) and the release year as the features of the movies. They are encoded as a $90$-dimensional multi-hot vector. We concatenate $Cy\in\{0,C\}$ to the feature vector to define the discrete distribution, where $y\in\{0,1\}$ indicates thumbs up or thumbs down, and $C=10$ is the hyperparameter that controls the emphasis on the label when we measure the distance between two data points. We focus on user 308, who is the first user who has $10$ thumbs up and $10$ thumbs down, and user 21, who is the first user who has more than $100$ ratings. For each user $u$, we aim to transfer the preference $\mathcal{D}_{u}\cap\mathcal{D}_{S}$ by choosing $K$ items from $\mathcal{D}_{T}$. We use MMD with the Gaussian kernel and bandwidth $\sigma=1$ as the distance measure. We use $L=1000$ iterations of the Frank-Wolfe algorithm and $R=100$ trials of the rounding process. We set $K=1,2,4,8,\ldots,128$ to see the convergence behavior.
@@ 6.2 Quantitative Evaluation
Datasets. We evaluate our method using datasets from three domains: MovieLens 100k (movies), Last.fm (music) [7], and Amazon-Home-and-Kitchen (e-commerce) [20, 32]. We use the same setting as the convergence analysis for the MovieLens 100K dataset except that we use all users. For the Last.fm dataset, we consider musicians with positive weights as thumbs up and we sample the same number of random musicians for thumbs down. We use the tags (e.g., metal, 80’s, chillout) as the features of the artists. As the number of tags is large, we apply the principal component analysis (PCA) to reduce the dimension to $50$ and standardize the variance. For the Amazon-Home-and-Kitchen dataset, we consider ratings $\geq 4$ as thumbs up and ratings <4<4 as thumbs down. We use the bag-of-words representation of the product reviews. We apply PCA to reduce the dimension to $50$ and standardize the variance. We consider two different settings for defining source and target services. The first setting, with intersection, is identical to the one used in the convergence analysis. The second setting, no intersection, is that we split the items into source and target services without any intersection. This setting is more challenging because the methods cannot copy any source items to the target service. For Last.fm and Amazon-Home-and-Kitchen datasets, we use the Gaussian kernel with the bandwidth $\sigma=10$ to reflect the scale of the features. We set $K=100$ for all datasets.
Baselines. Since there are no existing methods that solve the cold start problem by an end user, we use two simple baselines to validate that the proposed method is reasonably effective. The first baseline is random selection, which randomly selects $K$ items from $\mathcal{D}_{T}$. The second baseline is the greedy selection, which selects $K$ items in $J_{T}$ that have the smallest distance $d(x,\mathcal{D}_{S})=\min_{x^{\prime}\in\mathcal{D}_{S}}\|x-x^{\prime}\|_{2}$ to the source items $\mathcal{D}_{S}$. Additionally, we report the performance of the optimal solution of the continuous optimization problem.
Results. Table 1 shows MMD values between the source preference and the output of the methods. Lower is better. Pretender is much better than the baseline methods and is close to the optimal continuous solution. Note that checking the closeness is useful in practice as well. The continuous solution is obtained as a byproduct of Pretender. After we obtain the final output, we can verify that the final solution is indeed a reasonable one just by comparing the two values. We observe that the greedy method is worse than the random method. We hypothesize that it is because the greedy method chooses items that are too similar to some source items and the diversity is lost. We also observe that Pretender is robust to the setting with no intersection. The performance is slightly worse than in the setting with intersection but is still reasonablly well.